OpenStudy (anonymous):

help but hold on 1 sec ill put up the question in 1 sec

OpenStudy (anonymous):

 A. 8 B. 3 C. 10 D. 2

OpenStudy (anonymous):

is it d ??

OpenStudy (anonymous):

hello

hartnn (hartnn):

yes :) 2 is correct!

OpenStudy (anonymous):

ok 3 more

OpenStudy (anonymous):

Evaluate the expression when a = 4 and b = 6. 8a – 4b + 5 A. 15 B. 13 C. 7 D. 19

hartnn (hartnn):

what did u get ?

OpenStudy (anonymous):

is it b

hartnn (hartnn):

yes, 13 is correct :)

OpenStudy (anonymous):

Which is a variable expression for the word phrase? the product of 2 and a number decreased by 6 A. 2 + n – 6 B. 6 – 2n C. 2n + 6 D. 2n – 6

OpenStudy (anonymous):

d

hartnn (hartnn):

if d is \(2(n-6)\) then yes

OpenStudy (anonymous):

none of them have parinthisys

OpenStudy (anonymous):

( _)

hartnn (hartnn):

ok...best choice will still be D

OpenStudy (anonymous):

Select the mathematical sentence that is true. A. 4 • (6 – 2) + 5 > 9 + 3 • 4 B. 4 • (6 – 2) + 5 < 9 + 3 • 4 C. 4 • (6 – 2) + 5 = 9 + 3 • 4

hartnn (hartnn):

what u got ?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

a

hartnn (hartnn):

nopes

hartnn (hartnn):

what you got for left side? what you got for right side?

OpenStudy (anonymous):

b sorry did it rong and 16 and 21

OpenStudy (anonymous):

at first i got 39 and 21

hartnn (hartnn):

21 is correct 16 is incorrect how 16?

OpenStudy (anonymous):

ohh didint see 5 so c

OpenStudy (pinksapphire):

Yes.

OpenStudy (pinksapphire):

21=21

OpenStudy (anonymous):

Find the solution set for the given inequality if the replacement set is {1, 2, 3, 4, 5, 6}. 3x – 5 ≥ 7 A. {1, 2, 3, 4} B. {4, 5, 6} C. {1, 2, 3} D. {5, 6}

hartnn (hartnn):

did u try this ?

OpenStudy (anonymous):

i dont get theas

hartnn (hartnn):

3x – 5 ≥ 7 add 5 on both sides

OpenStudy (pinksapphire):

Plug in the numbers to x and see if it is true...

OpenStudy (pinksapphire):

So which numbers will work?

OpenStudy (anonymous):

b

OpenStudy (pinksapphire):

Correct :)

OpenStudy (anonymous):

Determine which value is a solution to the open sentence. 6 + 2x = 24 A. 9 B. 15 C. 16 D. 6

OpenStudy (pinksapphire):

Solve for x...

OpenStudy (anonymous):

i dont get it

OpenStudy (pinksapphire):

6+2x=24 -6 -6 2x=18 2x/2=18/2 x=9

OpenStudy (pinksapphire):

All you need to do is to get x to be by itself.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which is a word phrase for the variable expression? 4(n – 2) A. two less than four times a number n B. four times the product of a number n and 2 C. four more than the difference of a number n and 2 D. four times the difference of a number n and 2

OpenStudy (anonymous):

d

OpenStudy (anonymous):

???

OpenStudy (anonymous):

a or d i think

OpenStudy (anonymous):

im thinking its a do

OpenStudy (pinksapphire):

IDK maybe D.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes D

OpenStudy (anonymous):

i got 100%

OpenStudy (pinksapphire):

:)

hartnn (hartnn):

Congratulations!

OpenStudy (anonymous):

becus of u 2

OpenStudy (anonymous):

and tnx

hartnn (hartnn):

welcome ^_^

OpenStudy (pinksapphire):

No problem :)